# arXiv:2603.03190v1[cs.AI] 03 Mar 2026

### v1

@@ 1 Introduction
Within the predictive coding framework, music perception is described as a dynamic process of prediction and updating, in which sensory input is continuously processed through comparison with prior expectations [Koelsch2019, Vuust2022], and manipulation of such expectations can elicit musical pleasure and emotional responses [Huron2006, Juslin2008, Salimpoor2015]. A key element of musical expectation is melody; based on its statistical regularities, musical syntax is formed, and human listeners generate predictions about melodic continuation [Krumhansl2015, Patel2003, Rohrmeier2011, Tillmann2014]. Based on the free-energy principle [Friston2010], Friston positions music as “a domain in which the cognitive process of prediction is expressed in its purest form” [Friston2013]. Music perception is reformulated as a process of minimizing prediction error, and both perception and action are unified as inferential processes that minimize free energy, an upper bound on sensory surprisal [Friston2013]. For example, in rhythm perception, violations of predictions about beats and metrical structure give rise to prediction errors, and actions such as tapping modulate neural responses by adjusting predictive precision, consistent with active inference [Ishida2025].
@@
Listeners form both first-order predictions of perceptual content and second-order predictions of precision [Koelsch2019, Vuust2022]. From an information-theoretic perspective, expectation features can be described by Surprisal, reflecting the unexpectedness of realized events, and Entropy, quantifying uncertainty in the predictive distribution prior to events [Pearce2010]. These quantities are complementary and have been associated with separable neural response profiles [DiLiberto2020]. Accordingly, we describe expectation features along the two axes of Surprisal and Entropy.
@@
Auditory event-related potentials (ERPs) have been widely used as a methodological tool for investigating hierarchical expectation processing in the brain [Friston2013, Koelsch2011, Koelsch2008, Yu2015]. With respect to surprisal-related processing, several ERP components have been reported as neural indices of violations of musical expectations, including mismatch negativity (MMN) [Friston2013, Koelsch2011, Koelsch2008, Yu2015, Brattico2006, Mencke2021, QuirogaMartinez2020], early right anterior negativity (ERAN) [Friston2013, Koelsch2011, Koelsch2008, KoelschPLosONE], the N5 component [Friston2013, Koelsch2011, Koelsch2008, KoelschPLosONE], the N1 component [Carrus2013, QuirogaMartinez2020], the P300 response [Koelsch2008, QuirogaMartinez2020, Steinbeis2006], the P2 component [Koelsch2008, OMIGIE20131749, Choi2014, Heacock2019], late negative activity emerging around 400 ms [Pearce2010, MIRANDA2007331], and oscillatory activity [Pearce2010, Omigie2019]. In contrast, predictive uncertainty has been associated with decreases in MMN amplitude as a function of increasing Shannon entropy [Lumaca2019], and the stimulus-preceding negativity (SPN) has been reported as an index of uncertainty processing [Ono2024, Tanovic2019]. Importantly, although these ERP findings provide neural correlates of surprisal-like and entropy-like processes, most evidence has been obtained from experimental paradigms using artificial deviants rather than continuous naturalistic music [Koelsch2011, Koelsch2008].
@@
Conventional ERP-based paradigms rely on repeated stimuli and artificial deviants, limiting direct assessment of neural activity during naturalistic music listening. To address these limitations, encoding-model approaches have been applied to continuous naturalistic music using time-resolved regression frameworks such as the Temporal Response Function (TRF) [DiLiberto2020, Kern2023]. Within the lineage of encoding-model-based neural regression studies, expectation representations have been shown to be statistically dissociable from acoustic and adaptation effects [Kern2023, DiLiberto2020, GaleanoOtalaro2024], while acoustic features and ANN representations derived from audio or MIDI have likewise been demonstrated to be encoded in the human cortex [Mischler2025, Bellier2023, Tuckute2023].
@@
By leveraging evidence that acoustic-related ANN representations are encoded in the human cortex, Akama et al. proposed the PredANN framework, which im- proves EEG-based recognition [Akama2025]. EEG recognition plays a critical role in applications such as brain–computer interfaces (BCI) and neural decoding [Akama2025, Daly2023, Schirrmeister2017, Defossez2023]. For clarity, we refer to this framework hereafter as the PredANN framework. The PredANN framework predicts ANN representations from EEG, thereby compensating for information degradation and improving downstream recognition performance [Akama2025]. However, in the original PredANN framework, the supervisory ANN representations primarily captured acoustic structure, and the impact of expectation-related teacher representations on EEG recognition performance was not examined. Consequently, it cannot be concluded that acoustic features are optimal, nor was the potential complementarity between acoustic and expectation representations—suggested by neural regression studies—systematically tested.
@@
Here, inheriting the core hypothesis of the PredANN framework, we systematically examine how the choice of teacher representation itself influences EEG-based music identification and whether complementary effects emerge across distinct teacher types. Specifically, we pretrain models using ANN representations computed directly from music stimuli, explicitly distinguishing acoustic representations that predominantly encode acoustic information from expectation-related representations operationalized as Surprisal and Entropy. Fig. 1 illustrates the conceptual framework of our approach. We show that each representation independently contributes to performance improvements. Furthermore, integrating these representations yields complementary gains that surpass those achieved by simple but strong ensembles constructed solely through random initialization. The fact that recognition performance changes substantially depending on the context length used to compute expectation representations suggests that it is related to the context length of expectation processing in human music cognition. Peak recognition performance at relatively short context lengths is consistent with prior neuroscientific studies on music cognition [Kern2023]. Together, these results establish a new framework for designing EEG recognition models grounded in the intrinsic neural encoding structure of acoustic and expectation-related information during music perception. By demonstrating that ensemble diversity can be constructed through neurobiologically distinct representations rather than initialization differences, we show that representation learning design can be guided by the organization of information encoded in cortex, thereby redefining EEG model design on a neuroscientific foundation. This framework employs expectation features computed directly from raw signals, thereby enabling analysis of multilayer predictive structure in cortex that is not limited to onset or pitch. Directly computed from raw signals, these expectation features do not rely on symbolic representations such as MIDI or manual labels and can be naturally extended to diverse auditory stimuli. Extension to diverse auditory stimuli makes it possible to train models on diverse data and is therefore well aligned with the direction of foundation-style EEG models capable of solving various EEG tasks. This study contributes to improved performance and model interpretability in brain–computer interfaces and neural decoding technologies, and has the potential to advance understanding of predictive music cognition.
@@ 2 Results
Experiments were conducted using the NMED-T dataset [losorelli2017nmed], comprising EEG recordings from 20 participants listening to 10 distinct full-length musical pieces. The task is formulated as a 10-way identification problem, where each EEG segment is classified into its corresponding song ID; consequently, the chance-level accuracy is 0.1.
@@
Guided by predictive coding framework [Akama2025] and evidence for dissociable cortical encoding of acoustic and predictive information [DiLiberto2020], we asked a question: how does the choice of target representation affect downstream (Song ID classification) performance? We compared three neurophysiologically motivated representation types as pretraining targets: Acoustic, Surprisal, and Entropy. These representations reflect complementary aspects of auditory processing—acoustic properties of the signal versus predictive information about unexpected events (Surprisal) and uncertainty (Entropy). Acoustic features were extracted using MuQ [zhu2025muqselfsupervisedmusicrepresentation], and predictive features (Surprisal/Entropy) were computed using MusicGen [musicgen] (see Methods for details). We pretrained separate EEG encoders to predict each representation type, then fine-tuned them for Song ID classification. All three representation types were evaluated under identical experimental conditions (Seed 42, 16-s context window for predictive features). We refer to this framework, which leverages these complementary representation types as masked prediction targets, as PredANN++ (full architectural details are provided in Methods). This setup enables the direct comparison of how representation choice impacts EEG-based decoding performance (Figure 2).
@@
Di Liberto et al. have demonstrated that the human auditory cortex encodes predictive information, including surprise about unexpected auditory events and uncertainty about future predictions [DiLiberto2020]. These predictive aspects can be quantified using information-theoretic measures—Surprisal (quantifying unexpectedness) and Entropy (quantifying uncertainty)—computed from autoregressive models applied to the audio signal. The effectiveness of these predictive features for representation learning may depend on the temporal context length used to compute them, reflecting the temporal scope of cortical predictions during music listening.
To examine this context-dependency, we computed Surprisal and Entropy from an autoregressive music language model (MusicGen [musicgen]) using three context windows: 8 s, 16 s, and 32 s. Both Surprisal and Entropy demonstrated peaked accuracy at 16 s context (Figure 4). Surprisal at 16 s achieved 0.855 accuracy, improving over the full-scratch baseline by +3.2 pp, with all pairwise comparisons against individual seeds reaching significance by McNemar’s test ($p<0.001$). Entropy at 16 s attained 0.850 accuracy (+2.7 pp improvement), also with all seed-wise comparisons significant ($p<0.001$). In contrast, the 8-s and 32-s context windows showed more inconsistent performance, failing to achieve significant improvements across all seed baselines. This pattern suggests that the 16-s context length optimally captures the temporal structure of cortical predictions during music listening. We therefore adopted the 16 s configuration for all subsequent ensemble experiments.
@@
We additionally evaluated a conservative chunk-based scheme in which MusicGen logits are computed independently within each 30-s audio segment, and Surprisal and Entropy are derived accordingly (see Methods; results are reported in Supplementary Figure 1.).
@@
Prior studies have established that the human brain encodes both acoustic features [Bellier2023, Tuckute2023, Mischler2025] and expectation features [DiLiberto2020, Kern2023, GaleanoOtalaro2024] during music listening. Notably, the inclusion of melodic expectation features improves the prediction of neural responses beyond what is achieved by acoustic features alone [DiLiberto2020]. Based on this potential for information complementarity, we investigated whether integrating models trained with these distinct teacher signals—Surprisal, Entropy, and Acoustics—enhances EEG-based song ID classification. Specifically, we evaluated whether ensembling these models through probability averaging yields synergistic gains over any single-feature model.
@@
While the benefits of seed ensembling have been demonstrated in general computer vision tasks [DeepEnsemble2017, DeepEnsembleTheory2019], its effectiveness in the context of EEG-based song ID classification remains unexplored. Our results confirm that seed-based ensembling yields substantial gains in this domain (Figure 6). These robust improvements allow us to establish seed-based ensembling as a simple yet powerful baseline, against which the benefits of representation-based ensembling can be assessed in the subsequent section.
@@ 3 Discussion
ERP paradigms provide neurophysiological markers of expectation processing [Friston2013, Koelsch2011, Koelsch2008, Yu2015]. However, because they rely on artificially segmented stimulus sequences, these paradigms compromise stimulus naturalness and thereby constrain the study of natural music listening [DiLiberto2020]. Di Liberto et al. introduced TRF modelling to partially restore stimulus naturalness. However, their expectation features (Surprisal/Entropy) were derived from monophonic MIDI representations limited to pitch and onset time. As a result, these features remained restricted to onset-locked melodic statistics. Consequently, they did not define multidimensional expectation structures encompassing statistically predictable variations in duration, dynamics, amplitude envelope, timbral evolution, harmonic texture, instrumentation, reverberation, spectral structure, and other continuous acoustic attributes beyond discrete MIDI symbols.
@@
By contrast, in the present study we model predictive structure without relying on symbolic representations or explicit event segmentation. Specifically, we convert raw audio into EnCodec discrete tokens and compute Surprisal and Entropy from the next-token probabilities of MusicGen’s autoregressive distribution over these waveform-derived tokens. Because these next-token probabilities are defined directly on a representation trained to reconstruct the full audio signal, this formulation is holistic and multidimensional, capturing predictive regularities in natural music beyond pitch- and onset-level statistics [encodec, musicgen]. In what follows, we interpret the improvement in EEG-based Song ID classification accuracy reported in the Results section and discuss whether our framework offers a conceptually grounded modelling strategy for representation learning, including its neuroscientific motivation, limitations, and future directions.
@@
PredANN++ builds on Akama et al.’s hypothesis that predicting ANN representations from EEG provides an effective supervisory signal [Akama2025]. Unlike the supervised ANN representations used in Akama et al., we employ self-supervised learning models, namely MuQ [zhu2025muqselfsupervisedmusicrepresentation] and MusicGen [musicgen], consistent with evidence that self-supervised representations align more closely with cortical representations than supervised ones [NEURIPS2022_d81ecfc8, vaidya2022selfsupervisedmodelsaudioeffectively]. We use three teacher signals: an acoustic teacher derived from MuQ embeddings, and two expectation-related teachers derived from MusicGen, namely surprisal and entropy [Bellier2023, DiLiberto2020, Kern2023, Tuckute2023, GaleanoOtalaro2024, Mischler2025]. Surprisal quantifies the unexpectedness of an observed event, whereas entropy measures uncertainty in the predictive distribution prior to the event, and these quantities are distinguished as complementary information-theoretic measures [Pearce2010].
@@
Di Liberto et al. suggested that surprisal and entropy exhibit temporally complementary neural effects [DiLiberto2020]. However, in a neural regression study by Kern et al. (2023), surprisal effects were more readily detected than entropy effects, and adding entropy sometimes reduced model performance [Kern2023]. One possible explanation is that this discrepancy reflects methodological constraints inherent in onset-locked regression frameworks with correlated predictors, rather than directly implying that entropy lacks neural relevance. Uncertainty may manifest as gain-like modulation of deviance responses rather than as a strictly event-locked main effect [Lumaca2019]. Because surprisal is defined after an event and is therefore event-specific, whereas entropy reflects pre-event uncertainty in the predictive distribution, entropy may be difficult to capture using predictors computed only at note onsets. Moreover, surprisal and entropy are themselves correlated [DiLiberto2020], and incorporating correlated quantities within a single regression model can hinder reliable estimation of their unique contributions. Consistent with this concern, Galeano-Otálvaro et al. demonstrated entropy’s neural contribution by removing individual predictors from a full TRF model that included both acoustic and melodic predictors [GaleanoOtalaro2024]. For these reasons, we avoid forcing surprisal and entropy to compete within a single shared representation and instead pretrain separate encoders using each quantity as an independent teacher signal.
@@
Prior neural regression studies have shown that acoustic and melodic expectation features exhibit temporally and spatially dissociable neural responses. Temporally, acoustic-related responses typically emerge at earlier latencies (around $\sim 50$ ms), whereas expectation-related responses tend to arise at later time windows (around $\sim 200$ ms and beyond) [DiLiberto2020, Kern2023]. Spatially, acoustic responses are strongly associated with Heschl’s gyrus, while expectation-related activity extends to regions including the superior temporal gyrus and planum temporale [DiLiberto2020, Kern2023]. This spatiotemporal dissociation suggests that denoising aligned with Acoustic and Expectation can complement information across both time and cortical space. To test this directly, we trained a Transformer classifier to predict Song ID from Acoustic, Surprisal, or Entropy sequences alone, without EEG. In practice, we train an EEG encoder $F$ and decoder $G$ such that $G(F(x))$ approximates a stimulus-derived music feature $m$, thereby encouraging $F$ to learn representations predictive of $m$. A natural question is whether aligning EEG representations with $m$ benefits Song ID classification. To test this, we trained a Transformer classifier to predict Song ID from Acoustic, Surprisal, or Entropy sequences alone, without EEG. Supplementary Figure 2 shows that acoustic features, surprisal, and entropy each achieve near-perfect accuracy (approaching $100\%$) when used directly. Therefore, if these features are successfully recovered from EEG, they are sufficient for Song ID classification.
@@
We further adopt a deep-ensemble strategy by averaging class probabilities from encoders pretrained with different teacher signals. Deep ensembles improve performance when individual models produce less correlated errors [DeepEnsemble2017, DeepEnsembleTheory2019]. Unlike seed ensembles, which vary only random initialisation while keeping architecture, objective, and teacher fixed, our ensemble combines models pretrained with distinct teacher representations—acoustic features, surprisal, and entropy. These distinct teachers induce different inductive biases in the encoder $F$, encouraging more diverse error patterns and yielding gains beyond conventional seed-based ensembles.
@@
We next discuss the neuroscientific relevance of the context window used by MusicGen to compute Surprisal and Entropy. Kern et al. showed that EEG/MEG regression from a Music Transformer was most accurate when conditioning on fewer than roughly ten notes ($\sim$2–4 s) [Huang2019MusicTransformer]. In our study, varying the context window $W\in\{8,16,32\}$ revealed a clear peak at $W=16$. Although the tasks differ and direct comparison is not possible, the fact that performance is maximised at relatively short windows in both cases indicates a shared non-monotonic dependence on context length, supporting the neuroscientific plausibility of our formulation.
@@
Several factors may explain why a longer window is optimal here. First, unlike the monophonic MIDI-synthesised stimuli used by Kern et al., our stimuli are commercially produced polyphonic songs in which predictive structure may extend beyond ten-note contexts through rhythm, dynamics, acoustics, and energy. Second, unlike Liberto and Kern, who computed Surprisal only at discrete note-onset positions derived from MIDI, our Surprisal is computed at every 20-ms frame from a waveform-token language model. This frame-wise computation allows predictive quantities to be defined over the entire continuous time series rather than only at note events. As a result, our formulation can capture expectation-related structure that is not strictly tied to discrete note onsets, including slow dynamics and sub-note acoustic variations. A representation-learning objective that does not enforce strict onset locking may therefore better exploit these continuous predictive signals [Kern2023]. Third, because our Surprisal and Entropy are computed from a waveform-token language model (MusicGen over EnCodec tokens [musicgen, encodec]), they may reflect expectations not only for pitch/onset-like events but also for duration, dynamics, timbral evolution, and instrumentation changes unfolding over longer timescales. Importantly, although our encoder processes only a 3-s EEG segment, the teacher signals themselves are computed from longer context windows. By pretraining the model to predict these long-context quantities, the encoder is encouraged to infer predictive structure shaped by extended stimulus history from local neural activity. Given the representational capacity of Transformer architectures, such compressed long-range information can be embedded in the learned internal representation.

### v2

# Expectation and Acoustic Neural Network Representations Enhance Music Identification from Brain Activity

@@ 1 Introduction
Within the predictive coding framework, music perception is described as a dynamic process of prediction and updating, in which sensory input is continuously processed through comparison with prior expectations [1, 2], and manipulation of such expectations can elicit musical pleasure and emotional responses [3, 4, 5]. A key element of musical expectation is melody; based on its statistical regularities, musical syntax is formed, and human listeners generate predictions about melodic continuation [6, 7, 8, 9]. Based on the free-energy principle [10], Friston positions music as “a domain in which the cognitive process of prediction is expressed in its purest form” [11]. Music perception is reformulated as a process of minimizing prediction error, and both perception and action are unified as inferential processes that minimize free energy, an upper bound on sensory surprisal [11]. For example, in rhythm perception, violations of predictions about beats and metrical structure give rise to prediction errors, and actions such as tapping modulate neural responses by adjusting predictive precision, consistent with active inference [12].
@@
Listeners form both first-order predictions of perceptual content and second-order predictions of precision [1, 2]. From an information-theoretic perspective, expectation features can be described by Surprisal, reflecting the unexpectedness of realized events, and Entropy, quantifying uncertainty in the predictive distribution prior to events [13]. These quantities are complementary and have been associated with separable neural response profiles [14]. Accordingly, we describe expectation features along the two axes of Surprisal and Entropy.
@@
Auditory event-related potentials (ERPs) have been widely used as a methodological tool for investigating hierarchical expectation processing in the brain [11, 15, 16, 17]. With respect to surprisal-related processing, several ERP components have been reported as neural indices of violations of musical expectations, including mismatch negativity (MMN) [11, 15, 16, 17, 18, 19, 20], early right anterior negativity (ERAN) [11, 15, 16, 21], the N5 component [11, 15, 16, 21], the N1 component [22, 20], the P300 response [16, 20, 23], the P2 component [16, 24, 25, 26], late negative activity emerging around 400 ms [13, 27], and oscillatory activity [13, 28]. In contrast, predictive uncertainty has been associated with decreases in MMN amplitude as a function of increasing Shannon entropy [29], and the stimulus-preceding negativity (SPN) has been reported as an index of uncertainty processing [30, 31]. Importantly, although these ERP findings provide neural correlates of surprisal-like and entropy-like processes, most evidence has been obtained from experimental paradigms using artificial deviants rather than continuous naturalistic music [15, 16].
@@
Conventional ERP-based paradigms rely on repeated stimuli and artificial deviants, limiting direct assessment of neural activity during naturalistic music listening. To address these limitations, encoding-model approaches have been applied to continuous naturalistic music using time-resolved regression frameworks such as the Temporal Response Function (TRF) [14, 32]. Within the lineage of encoding-model-based neural regression studies, expectation representations have been shown to be statistically dissociable from acoustic and adaptation effects [32, 14, 33], while acoustic features and ANN representations derived from audio or MIDI have likewise been demonstrated to be encoded in the human cortex [34, 35, 36].
@@
By leveraging evidence that acoustic-related ANN representations are encoded in the human cortex, Akama et al. proposed the PredANN framework, which improves EEG-based recognition [37]. EEG recognition plays a critical role in applications such as brain–computer interfaces (BCI) and neural decoding [37, 38, 39, 40]. For clarity, we refer to this framework hereafter as the PredANN framework. The PredANN framework predicts ANN representations from EEG, thereby compensating for information degradation and improving downstream recognition performance [37]. However, in the original PredANN framework, the supervisory ANN representations primarily captured acoustic structure, and the impact of expectation-related teacher representations on EEG recognition performance was not examined. Consequently, it cannot be concluded that acoustic features are optimal, nor was the potential complementarity between acoustic and expectation representations—suggested by neural regression studies—systematically tested.
@@
Here, inheriting the core hypothesis of the PredANN framework, we systematically examine how the choice of teacher representation itself influences EEG-based music identification and whether complementary effects emerge across distinct teacher types. Specifically, we pretrain models using ANN representations computed directly from music stimuli, explicitly distinguishing acoustic representations that predominantly encode acoustic information from expectation-related representations operationalized as Surprisal and Entropy. Fig. 1 illustrates the conceptual framework of our approach. We show that each representation independently contributes to performance improvements. Furthermore, integrating these representations yields complementary gains that surpass those achieved by simple but strong ensembles constructed solely through random initialization. The fact that recognition performance changes substantially depending on the context length used to compute expectation representations suggests that it is related to the context length of expectation processing in human music cognition. Peak recognition performance at relatively short context lengths is consistent with prior neuroscientific studies on music cognition [32]. Together, these results establish a new framework for designing EEG recognition models grounded in the intrinsic neural encoding structure of acoustic and expectation-related information during music perception. By demonstrating that ensemble diversity can be constructed through neurobiologically distinct representations rather than initialization differences, we show that representation learning design can be guided by the organization of information encoded in cortex, thereby redefining EEG model design on a neuroscientific foundation. This framework employs expectation features computed directly from raw signals, thereby enabling analysis of multilayer predictive structure in cortex that is not limited to onset or pitch. Directly computed from raw signals, these expectation features do not rely on symbolic representations such as MIDI or manual labels and can be naturally extended to diverse auditory stimuli. Extension to diverse auditory stimuli makes it possible to train models on diverse data and is therefore well aligned with the direction of foundation-style EEG models capable of solving various EEG tasks. This study contributes to improved performance and model interpretability in brain–computer interfaces and neural decoding technologies, and has the potential to advance understanding of predictive music cognition.
@@ 2 Results
Experiments were conducted using the NMED-T dataset [41], comprising EEG recordings from 20 participants listening to 10 distinct full-length musical pieces. The task is formulated as a 10-way identification problem, where each EEG segment is classified into its corresponding song ID; consequently, the chance-level accuracy is 0.1.
@@
Guided by predictive coding framework [37] and evidence for dissociable cortical encoding of acoustic and predictive information [14], we asked a question: how does the choice of target representation affect downstream (Song ID classification) performance? We compared three neurophysiologically motivated representation types as pretraining targets: Acoustic, Surprisal, and Entropy. These representations reflect complementary aspects of auditory processing—acoustic properties of the signal versus predictive information about unexpected events (Surprisal) and uncertainty (Entropy). Acoustic features were extracted using MuQ [42], and predictive features (Surprisal/Entropy) were computed using MusicGen [43] (see Methods for details). We pretrained separate EEG encoders to predict each representation type, then fine-tuned them for Song ID classification. All three representation types were evaluated under identical experimental conditions (Seed 42, 16-s context window for predictive features). We refer to this framework, which leverages these complementary representation types as masked prediction targets, as PredANN++ (full architectural details are provided in Methods). This setup enables the direct comparison of how representation choice impacts EEG-based decoding performance (Figure 2).
@@
Di Liberto et al. have demonstrated that the human auditory cortex encodes predictive information, including surprise about unexpected auditory events and uncertainty about future predictions [14]. These predictive aspects can be quantified using information-theoretic measures—Surprisal (quantifying unexpectedness) and Entropy (quantifying uncertainty)—computed from autoregressive models applied to the audio signal. The effectiveness of these predictive features for representation learning may depend on the temporal context length used to compute them, reflecting the temporal scope of cortical predictions during music listening.
To examine this context-dependency, we computed Surprisal and Entropy from an autoregressive music language model (MusicGen [43]) using three context windows: 8 s, 16 s, and 32 s. Quantitative evaluation revealed that both Surprisal and Entropy demonstrated peaked accuracy at 16 s context (Figure 4). Surprisal at 16 s achieved 0.855 accuracy, improving over the full-scratch baseline by +3.2 pp, with all pairwise comparisons against individual seeds reaching significance by McNemar’s test ($p<0.001$). Entropy at 16 s attained 0.850 accuracy (+2.7 pp improvement), also with all seed-wise comparisons significant ($p<0.001$). In contrast, the 8-s and 32-s context windows showed more inconsistent performance, failing to achieve significant improvements across all seed baselines. Qualitative comparisons among the audio signals, the predictive features across the three context lengths, and the EEG recordings are also provided in Figure  12 to illustrate their temporal relationships (see also Supplementary Note 3: Qualitative inspection of context-dependent Surprisal/Entropy time courses for a detailed qualitative analysis). This overall pattern suggests that the 16-s context length optimally captures the temporal structure of cortical predictions during music listening. We therefore adopted the 16-s configuration for all subsequent ensemble experiments.
@@
We additionally evaluated a conservative chunk-based scheme in which MusicGen logits are computed independently within each 30-s audio segment, and Surprisal and Entropy are derived accordingly (see Methods; results are reported in Figure 10).
@@
Prior studies have established that the human brain encodes both acoustic features [35, 36, 34] and expectation features [14, 32, 33] during music listening. Notably, the inclusion of melodic expectation features improves the prediction of neural responses beyond what is achieved by acoustic features alone [14]. Based on this potential for information complementarity, we investigated whether integrating models trained with these distinct teacher signals—Surprisal, Entropy, and Acoustics—enhances EEG-based song ID classification. Specifically, we evaluated whether ensembling these models through probability averaging yields synergistic gains over any single-feature model.
@@
While the benefits of seed ensembling have been demonstrated in general computer vision tasks [44, 45], its effectiveness in the context of EEG-based song ID classification remains unexplored. Our results confirm that seed-based ensembling yields substantial gains in this domain (Figure 6). These robust improvements allow us to establish seed-based ensembling as a simple yet powerful baseline, against which the benefits of representation-based ensembling can be assessed in the subsequent section.
@@ 3 Discussion
ERP paradigms provide neurophysiological markers of expectation processing [11, 15, 16, 17]. However, because they rely on artificially segmented stimulus sequences, these paradigms compromise stimulus naturalness and thereby constrain the study of natural music listening [14]. Di Liberto et al. introduced TRF modelling to partially restore stimulus naturalness. However, their expectation features (Surprisal/Entropy) were derived from monophonic MIDI representations limited to pitch and onset time. As a result, these features remained restricted to onset-locked melodic statistics. Consequently, they did not define multidimensional expectation structures encompassing statistically predictable variations in duration, dynamics, amplitude envelope, timbral evolution, harmonic texture, instrumentation, reverberation, spectral structure, and other continuous acoustic attributes beyond discrete MIDI symbols.
@@
By contrast, in the present study we model predictive structure without relying on symbolic representations or explicit event segmentation. Specifically, we convert raw audio into EnCodec discrete tokens and compute Surprisal and Entropy from the next-token probabilities of MusicGen’s autoregressive distribution over these waveform-derived tokens. Because these next-token probabilities are defined directly on a representation trained to reconstruct the full audio signal, this formulation is holistic and multidimensional, capturing predictive regularities in natural music beyond pitch- and onset-level statistics [46, 43]. To allow readers to verify how these calculated metrics correspond to actual musical transitions, we provide an interactive web-based visualization at https://shogonoguchi.github.io/PredANNpp/#syncviz. The audio examples in this demonstration are drawn from the MTG-Jamendo dataset [47]. In what follows, we interpret the improvement in EEG-based Song ID classification accuracy reported in the Results section and discuss whether our framework offers a conceptually grounded modelling strategy for representation learning, including its neuroscientific motivation, limitations, and future directions.
@@
PredANN++ builds on Akama et al.’s hypothesis that predicting ANN representations from EEG provides an effective supervisory signal [37]. Unlike the supervised ANN representations used in Akama et al., we employ self-supervised learning models, namely MuQ [42] and MusicGen [43], consistent with evidence that self-supervised representations align more closely with cortical representations than supervised ones [48, 49]. We use three teacher signals: an acoustic teacher derived from MuQ embeddings, and two expectation-related teachers derived from MusicGen, namely surprisal and entropy [35, 14, 32, 36, 33, 34]. Surprisal quantifies the unexpectedness of an observed event, whereas entropy measures uncertainty in the predictive distribution prior to the event, and these quantities are distinguished as complementary information-theoretic measures [13].
@@
Di Liberto et al. suggested that surprisal and entropy exhibit temporally complementary neural effects [14]. However, in a neural regression study by Kern et al., surprisal effects were more readily detected than entropy effects, and adding entropy sometimes reduced model performance [32]. One possible explanation is that this discrepancy reflects methodological constraints inherent in onset-locked regression frameworks with correlated predictors, rather than directly implying that entropy lacks neural relevance. Uncertainty may manifest as gain-like modulation of deviance responses rather than as a strictly event-locked main effect [29]. Because surprisal is defined after an event and is therefore event-specific, whereas entropy reflects pre-event uncertainty in the predictive distribution, entropy may be difficult to capture using predictors computed only at note onsets. Moreover, surprisal and entropy are themselves correlated [14], and incorporating correlated quantities within a single regression model can hinder reliable estimation of their unique contributions. Consistent with this concern, Galeano-Otálvaro et al. demonstrated entropy’s neural contribution by removing individual predictors from a full TRF model that included both acoustic and melodic predictors [33]. For these reasons, we avoid forcing surprisal and entropy to compete within a single shared representation and instead pretrain separate encoders using each quantity as an independent teacher signal.
@@
Prior neural regression studies have shown that acoustic and melodic expectation features exhibit temporally and spatially dissociable neural responses. Temporally, acoustic-related responses typically emerge at earlier latencies (around $\sim 50$ ms), whereas expectation-related responses tend to arise at later time windows (around $\sim 200$ ms and beyond) [14, 32]. Spatially, acoustic responses are strongly associated with Heschl’s gyrus, while expectation-related activity extends to regions including the superior temporal gyrus and planum temporale [14, 32]. This spatiotemporal dissociation suggests that denoising aligned with Acoustic and Expectation can complement information across both time and cortical space. To test this directly, we trained a Transformer classifier to predict Song ID from Acoustic, Surprisal, or Entropy sequences alone, without EEG. In practice, we train an EEG encoder $F$ and decoder $G$ such that $G(F(x))$ approximates a stimulus-derived music feature $m$, thereby encouraging $F$ to learn representations predictive of $m$. A natural question is whether aligning EEG representations with $m$ benefits Song ID classification. To test this, we trained a Transformer classifier to predict Song ID from Acoustic, Surprisal, or Entropy sequences alone, without EEG. Figure 11 shows that acoustic features, surprisal, and entropy each achieve near-perfect accuracy (approaching $100\%$) when used directly. Therefore, if these features are successfully recovered from EEG, they are sufficient for Song ID classification.
@@
We further adopt a deep-ensemble strategy by averaging class probabilities from encoders pretrained with different teacher signals. Deep ensembles improve performance when individual models produce less correlated errors [44, 45]. Unlike seed ensembles, which vary only random initialisation while keeping architecture, objective, and teacher fixed, our ensemble combines models pretrained with distinct teacher representations—acoustic features, surprisal, and entropy. These distinct teachers induce different inductive biases in the encoder $F$, encouraging more diverse error patterns and yielding gains beyond conventional seed-based ensembles.
@@
We next discuss the neuroscientific relevance of the context window used by MusicGen to compute Surprisal and Entropy. Kern et al. showed that EEG/MEG regression from a Music Transformer was most accurate when conditioning on fewer than roughly ten notes ($\sim$2–4 s) [50]. In our study, varying the context window $W\in\{8,16,32\}$ revealed a clear peak at $W=16$. Although the tasks differ and direct comparison is not possible, the fact that performance is maximised at relatively short windows in both cases indicates a shared non-monotonic dependence on context length, supporting the neuroscientific plausibility of our formulation.
@@
Several factors may explain why a longer window is optimal here. First, unlike the monophonic MIDI-synthesised stimuli used by Kern et al., our stimuli are commercially produced polyphonic songs in which predictive structure may extend beyond ten-note contexts through rhythm, dynamics, acoustics, and energy. Second, unlike Liberto and Kern, who computed Surprisal only at discrete note-onset positions derived from MIDI, our Surprisal is computed at every 20-ms frame from a waveform-token language model. This frame-wise computation allows predictive quantities to be defined over the entire continuous time series rather than only at note events. As a result, our formulation can capture expectation-related structure that is not strictly tied to discrete note onsets, including slow dynamics and sub-note acoustic variations. A representation-learning objective that does not enforce strict onset locking may therefore better exploit these continuous predictive signals [32]. Third, because our Surprisal and Entropy are computed from a waveform-token language model (MusicGen over EnCodec tokens [43, 46]), they may reflect expectations not only for pitch/onset-like events but also for duration, dynamics, timbral evolution, and instrumentation changes unfolding over longer timescales. Importantly, although our encoder processes only a 3-s EEG segment, the teacher signals themselves are computed from longer context windows. By pretraining the model to predict these long-context quantities, the encoder is encouraged to infer predictive structure shaped by extended stimulus history from local neural activity. Given the representational capacity of Transformer architectures, such compressed long-range information can be embedded in the learned internal representation.